# Rethinking Table Instruction Tuning

### v2

@@ Abstract
Recent advances in table understanding have focused on instruction-tuning large language models (LLMs) for table-related tasks. However, existing research has overlooked the impact of hyperparameter choices, and also lacks a comprehensive evaluation of the out-of-domain table understanding ability and the general capabilities of these table LLMs. In this paper, we evaluate these abilities in existing table LLMs, and find significant declines in both out-of-domain table understanding and general capabilities as compared to their base models. Through systematic analysis, we show that hyperparameters, such as learning rate, can significantly influence both table-specific and general capabilities. Contrary to the previous table instruction-tuning work, we demonstrate that smaller learning rates and fewer training instances can enhance table understanding while preserving general capabilities. Based on our findings, we introduce TAMA, a TAble LLM instruction-tuned from LLaMA 3.1 8B Instruct, which achieves performance on par with, or surpassing GPT-3.5 and GPT-4 on table tasks, while maintaining strong out-of-domain generalization and general capabilities. Our findings highlight the potential for reduced data annotation costs and more efficient model development through careful hyperparameter selection. We open-source the project and our models.111 Our project page is at https://lit.eecs.umich.edu/TAMA/. We open-source our code at https://github.com/MichiganNLP/TAMA. We open-source our models at https://huggingface.co/MichiganNLP/tama-5e-7 and https://huggingface.co/MichiganNLP/tama-1e-6. We open-source the data used in this paper at https://huggingface.co/datasets/MichiganNLP/TAMA_Instruct.
@@ 2.2 Findings
Existing Table LLMs demonstrate poor instruction-following ability. In Table 3, both TableLLaMA and TableLLM show significant drops in performance on IFEval (Zhou et al., 2023), with accuracy declines of 5.63 and 17.86, resulting in a score of 25.78 and 30.46, respectively. While TableBenchLLM maintains a similar score to its base model (32.85 compared to 32.13 for LLaMA 3.1-8B), this performance is still limited compared to 83.57 by GPT-4 reported by Zhou et al. (2023). At such low instruction following scores, existing table LLMs cannot consistently follow instructions such as “return the answer in JSON format” as shown in Table 7 in Section 4.3 and Tables 17, 18 and 19 in Appendix E, limiting the model’s usage if the end users need data extraction that requires certain answer format.
Existing table instruction tuning compromises models’ general capabilities. Existing table instruction-tuning methods lead to significant drops in accuracy on general benchmarks such as MMLU, AI2ARC, GPQA as shown in Table 3. For instance, compared to their base models, TableLLaMA experiences a decline of 13.95 accuracy score on MMLU, while TableLLM and TableBenchLLM lose 8.79 and 9.41, respectively. Appendix B provides further discussion of the model’s performance corresponding to each category in MMLU benchmark. On the general reasoning benchmarks such as AI2ARC, the drop can be as large as 20.90 for TableBenchLLM, showing that the existing table instruction tuning hurts their base model’s reasoning ability. This limits the existing table LLMs’ usage if there are general knowledge or reasoning involved in end users’ request.
@@ 3.1 Experimental Setup
We conduct full parameter table instruction tuning using the 8B version of the LLaMA 3.1 Instruct model (Dubey et al., 2024) because of its superior general capabilities, especially its strong instruction following ability. Section C.1 provides detailed reasons for the choice of LLaMA 3.1 and the choice between the base versus the instruction-tuned version of the model. In addition, we expand our analyses across various LLMs and learning setups detailed in Section 3.2.
@@ 3.2 Analysis
    Learning rate across LLMs under full parameter setup (Section C.4), and LoRA and QLoRA setups (Section C.5).
    Finding: We list our recommended learning rate per model in Table 4 for the full parameter setup. Please refer to Section C.5 for recommendations for LoRA and QLoRA setups.
@@
    Number of examples across LLMs under full parameter tuning setup (Section C.4), and LoRA and QLoRA setups (Section C.5).
@@
    Number of epochs (Section C.2).
@@
    Multi-task training (Section C.3).
@@
In addition, we provide analysis regarding how the data features affect the model’s performance degradation on general benchmarks in Section C.7.
@@ 4.1 Experimental Setup
Appendix A provides more details of our experimental setup including the information of GPU server, generation hyperparameters, data processing, and our evaluation setup. Appendix F provides examples from datasets that we evaluate upon.
@@ 4.2 Results and Analyses
Table 7: Table LLMs’ predictions on the prompts from IFEval and Table-Syn (S1 in Table 6). We omit parts of the examples for readability. Appendix E provides the complete examples.
@@ 4.3 Hindsight Analysis
In hindsight, we want to validate that our selected hyperparameters indeed work the best. Therefore, we run the experiments on the same training set with the learning rate ranging from 1.0e-7 to 1.0e-5, and the number of epochs from one to six. Figure 3 reports part of the results, and Appendix D reports the complete results and provide further discussion.
@@ 6 Conclusion
Due to the space constraint, we provide further analysis across LLMs in Section C.4, including Llama 2 7B Instruct (Touvron et al., 2023), QWen 2.5 7B Instruct (Bai et al., 2023), Mistral v0.3 7B Instruct (Jiang et al., 2023), and Phi 3 small 8K Instruct (7B) (Abdin et al., 2024), and analysis in terms of LoRA and QLoRA in Section C.5. We provide further analysis regarding how the data features affect the model’s performance degradation on general benchmarks in Section C.7. However, due to the scope of one study, we cannot exhaust every possible foundational model and every possible dataset.
@@ Acknowledgements
Appendix A Experiment Details
A.1 GPU Details
@@ A.3 Batch Sizes
A.4 Details of Prompting GPT Models
@@
A.5 Details of Data Processing
@@ A.6 Details of Evaluation
Appendix B Evaluation of the Existing Table LLMs.
@@
Appendix C Model and Hyperparameter Exploration
C.1 Model Selection
@@ C.2 Effects of Epochs
C.3 Effects of Multi-Task
In Figure 6, we present the heatmap of model performance when fine-tuning the LLaMA 3.1 8B Instruct model on a single dataset (one of the datasets among FeTaQA, HiTab, and TabFact). We fine-tune the model for two epochs at a learning rate of 1.0e-6 with 500 instruction pairs, and then test it against the six datasets. Additionally, Figures 7 and 8 in Section C.6 present heatmaps across varying learning rates (from 1.0e-7 to 1.0e-5) and number of epochs (from one to six).
@@ C.3 Effects of Multi-Task
C.4 Hyperparameter Exploration Across Models
@@
C.5 Hyperparameter Exploration for LoRA and QLoRA
@@
We replicate the experiments we conduct in Section C.4, and here we present our results in two aspects, the learning rate and the number of examples.
@@
C.6 Individual Task’s Influence on Model Performance
@@
C.7 Trade-off Analysis for Data Properties
@@
Appendix D Hindsight Analysis
@@
Appendix E Model Prediction Examples
Table 17 provides an example for table LLMs’ generation on IFEval dataset. Tables 18 and 19 provide two examples for table LLM’s generation on Table-Syn dataset. Apart from the limited out-of-domain table reasoning ability, we find that existing table LLMs also exhibit limited instruction-following capabilities, and often struggle with consistently returning answers in specified formats, such as JSON. Such a limitation poses challenges in the practical use cases, where the end-users may request specific output formats to extract answers from the model’s predictions.
@@ Appendix E Model Prediction Examples
Appendix F Dataset Examples
F.1 WikiTQ

### v3

@@ Abstract
Recent advances in table understanding have focused on instruction-tuning large language models (LLMs) for table-related tasks. However, existing research has overlooked the impact of hyperparameter choices, and also lacks a comprehensive evaluation of the out-of-domain table understanding ability and the general capabilities of these table LLMs. In this paper, we evaluate these abilities in existing table LLMs, and find significant declines in both out-of-domain table understanding and general capabilities as compared to their base models. Through systematic analysis, we show that hyperparameters, such as learning rate, can significantly influence both table-specific and general capabilities. Contrary to the previous table instruction-tuning work, we demonstrate that smaller learning rates and fewer training instances can enhance table understanding while preserving general capabilities. Based on our findings, we introduce TAMA, a TAble LLM instruction-tuned from LLaMA 3.1 8B Instruct, which achieves performance on par with, or surpassing GPT-3.5 and GPT-4 on table tasks, while maintaining strong out-of-domain generalization and general capabilities. Our findings highlight the potential for reduced data annotation costs and more efficient model development through careful hyperparameter selection. We open-source the project and our models.
Updates:
@@ 2.2 Findings
Existing Table LLMs demonstrate poor instruction-following ability. In Table 3, both TableLLaMA and TableLLM show significant drops in performance on IFEval (Zhou et al., 2023), with accuracy declines of 5.63 and 17.86, resulting in a score of 25.78 and 30.46, respectively. While TableBenchLLM maintains a similar score to its base model (32.85 compared to 32.13 for LLaMA 3.1-8B), this performance is still limited compared to 83.57 by GPT-4 reported by Zhou et al. (2023). At such low instruction following scores, existing table LLMs cannot consistently follow instructions such as “return the answer in JSON format” as shown in Table 7 in Section 4.3 and Tables 18, 19 and 20 in Appendix F, limiting the model’s usage if the end users need data extraction that requires certain answer format.
Existing table instruction tuning compromises models’ general capabilities. Existing table instruction-tuning methods lead to significant drops in accuracy on general benchmarks such as MMLU, AI2ARC, GPQA as shown in Table 3. For instance, compared to their base models, TableLLaMA experiences a decline of 13.95 accuracy score on MMLU, while TableLLM and TableBenchLLM lose 8.79 and 9.41, respectively. Appendix C provides further discussion of the model’s performance corresponding to each category in MMLU benchmark. On the general reasoning benchmarks such as AI2ARC, the drop can be as large as 20.90 for TableBenchLLM, showing that the existing table instruction tuning hurts their base model’s reasoning ability. This limits the existing table LLMs’ usage if there are general knowledge or reasoning involved in end users’ request.
@@ 3.1 Experimental Setup
We conduct full parameter table instruction tuning using the 8B version of the LLaMA 3.1 Instruct model (Dubey et al., 2024) because of its superior general capabilities, especially its strong instruction following ability. Section D.1 provides detailed reasons for the choice of LLaMA 3.1 and the choice between the base versus the instruction-tuned version of the model. In addition, we expand our analyses across various LLMs and learning setups detailed in Section 3.2.
@@ 3.2 Analysis
    Learning rate across LLMs under full parameter setup (Section D.4), and LoRA and QLoRA setups (Section D.5).
    Finding: We list our recommended learning rate per model in Table 4 for the full parameter setup. Please refer to Section D.5 for recommendations for LoRA and QLoRA setups.
@@
    Number of examples across LLMs under full parameter tuning setup (Section D.4), and LoRA and QLoRA setups (Section D.5).
@@
    Number of epochs (Section D.2).
@@
    Multi-task training (Section D.3).
@@
In addition, we provide analysis regarding how the data features affect the model’s performance degradation on general benchmarks in Section D.7.
@@ 4.1 Experimental Setup
Appendix B provides more details of our experimental setup including the information of GPU server, generation hyperparameters, data processing, and our evaluation setup. Appendix G provides examples from datasets that we evaluate upon.
@@ 4.2 Results and Analyses
Table 7: Table LLMs’ predictions on the prompts from IFEval and Table-Syn (S1 in Table 6). We omit parts of the examples for readability. Appendix F provides the complete examples.
@@ 4.3 Hindsight Analysis
In hindsight, we want to validate that our selected hyperparameters indeed work the best. Therefore, we run the experiments on the same training set with the learning rate ranging from 1.0e-7 to 1.0e-5, and the number of epochs from one to six. Figure 3 reports part of the results, and Appendix E reports the complete results and provide further discussion.
@@ 6 Conclusion
Due to the space constraint, we provide further analysis across LLMs in Section D.4, including Llama 2 7B Instruct (Touvron et al., 2023), QWen 2.5 7B Instruct (Bai et al., 2023), Mistral v0.3 7B Instruct (Jiang et al., 2023), and Phi 3 small 8K Instruct (7B) (Abdin et al., 2024), and analysis in terms of LoRA and QLoRA in Section D.5. We provide further analysis regarding how the data features affect the model’s performance degradation on general benchmarks in Section D.7. However, due to the scope of one study, we cannot exhaust every possible foundational model and every possible dataset.
@@ Acknowledgements
Appendix A Updates
A.1 Better TAMA Models with Limited Data
Overall
Models Training Corpora Size Base Model Model Size Performance
TableLLaMA (2024a) 2M LongLoRA 7B 0.0
TableLLM (2024b) 309K CodeLLaMA 7B 2.5
TableBenchLLM (2024) 20K LLaMA 3.1 8B 3.4
LLaMA 3.1 8B Instruct (2024) - LLaMA 3.1 8B 25.3
TAMA-vB (ours) 2.6K LLaMA 3.1 8B 21.1
TAMA-vA (ours) 2.6K LLaMA 3.1 8B 16.9
Qwen2.5 7B (2025) - QWen 2.5 7B 28.5
TableGPT2 (2024) 2.36M QWen 2.5 7B 30.0
Table R1 Zero (2025b) 48.6K LLaMA 3.1 8B 26.6
TAMA-QWen2.5 (ours) 2.6K QWen 2.5 7B 27.6
QWen 3 8B (2025a) - QWen 3 8B 32.9
TAMA-QWen3 (ours) 2.6K QWen 3 8B 33.9
Table 10: Performance comparisons on MMTU (Xing et al., 2025). TAMA series yield competitive performance with limited training data. Notably, TAMA-QWen3 yields the best performance on MMTU of 33.9 among 7B and 8B models.
In this paper, we reveal that with limited instruction tuning data, we can achieve competitive performance on table tasks. This compact setup enables quick instruction tuning with advanced base models.
We present TAMA models built on Qwen 2.5 and Qwen 3. These models achieve strong results on the MMTU benchmark (Xing et al., 2025), outperforming recent table reasoning models (Yang et al., 2025b) and competitive table LLMs like Table-GPT 2 (Su et al., 2024), which is tuned on 2.36M datapoints.
Notably, in Table 10, TAMA-QWen3 achieves the best overall performance of 33.9, surpassing QWen-3-8B (32.9) and TableGPT-2 (30.0).
We adopt the official MMTU evaluation script to compute scores. For QWen 3 model (Yang et al., 2025a) and TAMA-QWen3, we turned off the thinking mode.
Appendix B Experiment Details
B.1 GPU Details
@@ B.3 Batch Sizes
B.4 Details of Prompting GPT Models
@@
B.5 Details of Data Processing
@@ B.6 Details of Evaluation
Appendix C Evaluation of the Existing Table LLMs.
@@
Appendix D Model and Hyperparameter Exploration
D.1 Model Selection
@@ D.2 Effects of Epochs
D.3 Effects of Multi-Task
In Figure 6, we present the heatmap of model performance when fine-tuning the LLaMA 3.1 8B Instruct model on a single dataset (one of the datasets among FeTaQA, HiTab, and TabFact). We fine-tune the model for two epochs at a learning rate of 1.0e-6 with 500 instruction pairs, and then test it against the six datasets. Additionally, Figures 7 and 8 in Section D.6 present heatmaps across varying learning rates (from 1.0e-7 to 1.0e-5) and number of epochs (from one to six).
@@ D.3 Effects of Multi-Task
D.4 Hyperparameter Exploration Across Models
@@
D.5 Hyperparameter Exploration for LoRA and QLoRA
@@
We replicate the experiments we conduct in Section D.4, and here we present our results in two aspects, the learning rate and the number of examples.
@@
D.6 Individual Task’s Influence on Model Performance
@@
D.7 Trade-off Analysis for Data Properties
@@
Appendix E Hindsight Analysis
@@
Appendix F Model Prediction Examples
Table 18 provides an example for table LLMs’ generation on IFEval dataset. Tables 19 and 20 provide two examples for table LLM’s generation on Table-Syn dataset. Apart from the limited out-of-domain table reasoning ability, we find that existing table LLMs also exhibit limited instruction-following capabilities, and often struggle with consistently returning answers in specified formats, such as JSON. Such a limitation poses challenges in the practical use cases, where the end-users may request specific output formats to extract answers from the model’s predictions.
@@ Appendix F Model Prediction Examples
Appendix G Dataset Examples
G.1 WikiTQ